The volume of a cone varies jointly as its height and the square of its radius. A certain cone has a volume of 720 pi cm cubed , a height of 15 cm , and a radius of 12 cm. Find the radius of another cone that has a height of 27 cm and a volume of 567pi cm cubed
Answer by jorel1380(3719)

1/3πr2h=567
πr2h=1701
πr2=63
r2=63/π=21
r=√21..